ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 53 Aminova Asya V., Kazan State University, Russia The pseudo-Riemannian manifolds with common geodesics It is given the solution of the classic geometrical problem of determination of all pseudo-Riemannian metrics with corresponding geodesics that has appeared in connection with dynamic problem about transformations preserving trajectories and more than one hundred years later from the times of Beltrami, Dini and Levi-Civita whose theorems originated the problem it was still not solved. Reference: A. V. Aminova. Pseudo-Riemannian manifolds with common geodesics, Uspekhi Matematicheskikh Nauk, 48 (1993), iss. 2 (290), 107– 164. [Russian Math. Syrveys 48 (1993), No 2, 105–160. ]